What mass of carbon dioxide is formed by the complete combustion of 775 g of ethanol? CH3CH2OH + O2 => CO2 + H2O?
1
Expert's answer
2017-06-23T09:17:07-0400
2 CH3CH2OH + 7 O2 = 4 CO2 + 6H2O

n(ethanol) = m/M = 775 g / 47 g/mol = 16.85 mol
n(CO2) = n(ethanol) / 2 * 4 (according to the reaction coefficients) = 33.90 mol
m(CO2) = n*M = 44*33.9 = 1491.6 g